# Map It Anywhere (MIA): Empowering Bird’s Eye View Mapping using Large-scale Public Data

### v1

@@ Abstract
Top-down Bird’s Eye View (BEV) maps are a popular representation for ground robot navigation due to their richness and flexibility for downstream tasks. While recent methods have shown promise for predicting BEV maps from First-Person View (FPV) images, their generalizability is limited to small regions captured by current autonomous vehicle-based datasets. In this context, we show that a more scalable approach towards generalizable map prediction can be enabled by using two large-scale crowd-sourced mapping platforms, Mapillary for FPV images and OpenStreetMap for BEV semantic maps. We introduce Map It Anywhere (MIA), a data engine that enables seamless curation and modeling of labeled map prediction data from existing open-source map platforms. Using our MIA data engine, we display the ease of automatically collecting a dataset of 1.2 million pairs of FPV images & BEV maps encompassing diverse geographies, landscapes, environmental factors, camera models & capture scenarios. We further train a simple camera model-agnostic model on this data for BEV map prediction. Extensive evaluations using established benchmarks and our dataset show that the data curated by MIA enables effective pretraining for generalizable BEV map prediction, with zero-shot performance far exceeding baselines trained on existing datasets by 35%. Our analysis highlights the promise of using large-scale public maps for developing & testing generalizable BEV perception, paving the way for more robust autonomous navigation.
@@ 1 Introduction
Bird’s Eye View (BEV) maps are an important perceptual representation for ground robots. Across various applications such as autonomous driving [19, 3, 13, 29], offroad navigation [24, 33, 1, 17], and localization [30, 17], prior work has demonstrated the simplicity of reasoning through the BEV perspective, making it also attractive for many downstream tasks such as path planning [26, 14]. To enable wide deployment of BEV maps as a perceptual representation, there is a requirement for a general map prediction building block that performs robustly across different domains and supports effective adaptation to specific tasks/environments.
@@
    Using our MIA data engine, we release a dataset containing 1.2 million high quality FPV image and BEV map pairs covering 470 $km^{2}$, thereby facilitating future map prediction research on generalizability and robustness.
@@
Dataset Locations km2 covered # BEV Annotated Frames # Camera Models Domain types Capture Platform Automatic Curation
@@ 2 Related Work
Crowd-sourced Datasets for Learning Geometric Tasks: Crowd-sourced platforms enable open-source contributors to upload diverse in-the-wild data, significantly empowering generalizability in geometric learning tasks. One such notable platform is Mapillary [7], which hosts over 2 billion (and growing) crowd-sourced street-level images from various locations worldwide, captured by different cameras across all seasons and times. Mapillary has been notably used for tasks such as depth estimation [2], lifelong place recognition [35], and visual localization [16, 30]. Most related to our work, OrienterNet [30] addresses visual localization within a large top-down map, curating a large-scale localization dataset provided by crowd-sourced platforms, Mapillary [7] for FPV images and OpenStreetMap [8] (OSM) for large BEV maps. However, the OrienterNet pipeline for rasterizing maps is not suitable for BEV prediction as it renders large maps, mimics OSM style, and includes graphical elements/labels irrelevant to the map prediction task. MIA further refines the OrienterNet pipeline by enabling automatic curation and collection of semantic maps, alignment of BEV renders with satellite images, and inference of missing OSM sidewalk geometries, thus providing rich semantic maps that are ready for BEV map prediction.
@@ 3.1 First Person View (FPV) Retrieval
FPV Pipeline: As demonstrated in Fig. 2, the FPV pipeline starts by manually inputting a list of locations of interest, which can be as simple as inputting the name of the location or as specific as specifying the GPS bounds. The pipeline then converts these bounds to a list of zoom-14 tiles and uses the publicly exposed Mapillary APIs to query for any image instance within these tiles. The retrieved instances are then geospatially filtered to ensure they lie within the exact boundaries of interest. Given the retrieved image IDs, we use another Mapillary endpoint to retrieve image metadata, which includes coordinates rectified through structure from motion, camera information, poses, and timestamps, amongst other details that we use for filtering in the subsequent stage.
We develop the filtering pipeline by observing hundreds of FPV & BEV pairs and identifying the correlations between good-quality FPVs and their corresponding metadata. The criteria we used included a recency filter, a camera model filter spanning 19 camera models with good RGB quality, a location/angle discrepancy filter that computes the difference between Structure from Motion (SfM) computed and recorded poses as a proxy for measuring the quality of the geo-registration, and a camera type filter that only includes perspective and fisheye. To promote spatial diversity over sheer quantity, we filter out images within a 4-meter radius of another image from the same sequence. After filtering, we retrieve the RGB images from Mapillary and feed them through an undistortion pipeline adapted from [30]. The undistortion is critical for fisheye images to ensure their pixel-aligned features can be correctly lifted into BEV space. Using this pipeline, we can retrieve high-quality images from anywhere in the world, tapping into the power of the Mapillary platform.
@@ 7 Conclusion
In this work, we propose MIA, a data curation pipeline aimed at empowering anywhere BEV map prediction from FPV images. We release a large MIA dataset obtained through the data engine giving the research community access to 1.2M FPV-BEV pairs to accelerate anywhere map prediction research. Results from training on the dataset show impressive generalization performance across conventional map prediction datasets, while on the other hand, provides challenging test cases for the research community. Our approach departs from the traditional and expensive autonomous vehicle data collection and labeling paradigm, towards automatic curation of readily-available crowd-sourced data. We believe this work seeds the first step towards anywhere map prediction.
@@ 7.1 Limitations, Biases, Social Impact
While we show promising generalization performance on conventional datasets, we note that label noise inherently exists, to a higher degree than manually collected data, in crowd-sourced data, in both pose correspondence and in BEV map labeling. Such noise is common across large-scale automatically scraped/curated benchmarks such as ImageNet [10]. Moreover, our approach does not capture dynamic classes as they do not exist in static maps. However, we see our approach as indispensable for scale and diversity yet complementary to conventionally obtained datasets. While we recognize that our sampled dataset is biased towards locations in the US, our MIA data engine is directly applicable to other worldwide locations.
@@
We thank Mihir Sharma for his work on related research curation and exploration into potential extensions involving terrain height extraction. We thank Mononito Goswami for invaluable discussions on paper organization, framing, experimental design, and project feedback. Additionally, we appreciate David Fan, Simon Stepputtis, and Yaqi Xie for their insightful feedback throughout the project.
@@ A.3 Intra-Dataset Class Mappings
A.4 Hyperparameters
@@
A.5 Dataset Privacy and Consent
@@
Appendix B Additional MIA data samples
Figure 9: Additional MIA data samples and associated metadata
Appendix C Mapper Model Architecture
Figure 10: Mapper Architecture

### v2

@@ Abstract
Top-down Bird’s Eye View (BEV) maps are a popular perceptual representation for ground robot navigation due to their richness and flexibility for downstream tasks. While recent methods have shown promise for predicting BEV maps from First-Person View (FPV) images, their generalizability is limited to small regions captured by current autonomous vehicle-based datasets. In this context, we show that a more scalable approach towards generalizable map prediction can be enabled by using two large-scale crowd-sourced mapping platforms, Mapillary for FPV images and OpenStreetMap for BEV semantic maps. We introduce Map It Anywhere (MIA), a data engine that enables seamless curation and modeling of labeled map prediction data from existing open-source map platforms. Using our MIA data engine, we display the ease of automatically collecting a dataset of 1.2 million pairs of FPV images & BEV maps encompassing diverse geographies, landscapes, environmental factors, camera models & capture scenarios. We further train a simple camera model-agnostic model on this data for BEV map prediction. Extensive evaluations using established benchmarks and our dataset show that the data curated by MIA enables effective pretraining for generalizable BEV map prediction, with zero-shot performance far exceeding baselines trained on existing datasets by 35%. Our analysis highlights the promise of using large-scale public maps for developing & testing generalizable BEV perception, paving the way for more robust autonomous navigation.
@@ 1 Introduction
Bird’s Eye View (BEV) maps are an important perceptual representation for localization[33, 20], mapping [32, 15, 22, 3, 29], and decision-making tasks [27, 18], particularly for ground robots. They provide a rich, efficient metric representation of the world, enabling spatial reasoning directly in the plane ground robots move in. Given these advantages, autonomous driving systems often employ BEV maps as their primary perceptual representation [42, 18]. Beyond on-road driving, BEV mapping is widely used in other robot domains, such as offroad driving[27, 36, 1], mobile manipulation [40], and exploration [12, 16]. To enable wide deployment of BEV maps as a perceptual representation, there is a requirement for a general map prediction building block that performs robustly across different domains and supports effective adaptation to specific tasks/environments.
@@
    Using our MIA data engine, we release a dataset containing $\sim$1.2 million high quality FPV image and BEV map pairs covering 470 $km^{2}$, thereby facilitating future map prediction research on generalizability and robustness.
@@
Dataset Locations km2 covered # Annotated BEV Frames # Camera Models Domain types Capture Platform Suitable for Map Pred. Automatic Curation
@@ 2 Related Work
Crowd-sourced Datasets for Learning Geometric Tasks: Crowd-sourced platforms enable open-source contributors to upload diverse in-the-wild data, significantly empowering generalizability in geometric learning tasks. One such notable platform is Mapillary [7], which hosts over 2 billion (and growing) crowd-sourced street-level images from various locations worldwide, captured by different cameras across all seasons and times. Mapillary has been notably used for tasks such as depth estimation [2], lifelong place recognition [38], and visual localization [19, 33]. Most related to our work, OrienterNet [33] addresses visual localization within a large top-down map, curating a large-scale localization dataset provided by crowd-sourced platforms, Mapillary [7] for FPV images and OpenStreetMap [9] (OSM) for large BEV maps. However, the OrienterNet pipeline for rasterizing maps is not suitable for BEV prediction as it renders large maps, mimics OSM style, and includes graphical elements/labels irrelevant to the map prediction task. We illustrate the qualitative differences in Figs. 11 and 12 (shared in the Appendix). MIA further refines the OrienterNet pipeline by enabling automatic curation and collection of semantic maps, alignment of BEV renders with satellite images, and inference of missing OSM sidewalk geometries, thus providing rich semantic maps that are ready for BEV map prediction.
@@ 3.1 First Person View (FPV) Retrieval
FPV Pipeline: As demonstrated in Fig. 2, the FPV pipeline starts by manually inputting a list of locations of interest, which can be as simple as inputting the name of the location or as specific as specifying the GPS bounds. The geographical bounds are then fetched using the Nominatim API [8] if needed. The pipeline then converts these bounds to a list of zoom-14 tiles and uses the public Mapillary APIs to query for the image instances within these tiles. Only the retrieved instances that lie within the geographical boundaries of interest are kept. Given the retrieved image IDs, we use another Mapillary endpoint to retrieve image metadata, which includes coordinates rectified through structure from motion, camera information, poses, and timestamps, amongst other details that we use for filtering in the subsequent stage. For more information on the raw data used from Mapillary, please refer to Supplemental Information Q8.
We develop the filtering pipeline by observing hundreds of FPV & BEV pairs and identifying the correlations between good-quality FPVs and their corresponding metadata. The criteria we used included a recency filter, a camera model filter spanning 19 camera models with good RGB quality, a location/angle discrepancy filter that computes the difference between Structure from Motion (SfM) computed and recorded poses, both obtained from Mapillary API, as a proxy for measuring the quality of the geo-registration, and a camera type filter that only includes perspective and fisheye. To promote spatial diversity over sheer quantity, we filter out images within a 4-meter radius of another image from the same sequence. After filtering, we retrieve the RGB images from Mapillary and feed them through an undistortion pipeline adapted from [33]. The undistortion is critical for fisheye images to ensure their pixel-aligned features can be correctly lifted into BEV space. Using this pipeline, we can retrieve high-quality images from anywhere in the world, tapping into the power of the Mapillary platform.
@@ 6.3 MIA provides challenging settings for future work on anywhere map prediction
Figure 8: Example pairs of First-Person-View images & Birds-Eye-View semantic maps effortlessly extracted by our MIA data engine from non-US cities.
Figure 9: Zero-shot predictions in non-US cities. Mapper consistently provides more accurate & realistic zero-shot predictions.
Methods Avg. over Classes Avg. R, S
TIIM [32] 10.78 21.21
SkyEye [15] 13.77 20.87
Mapper 17.66 37.26
Table 3: Zero-shot IoU results on MIA-non-US data split.
6.4 MIA data engine can also effortlessly extract data beyond U.S. cities
To show that our MIA data engine can effortlessly curate data from other global locations, we curate MIA-non-US from 6 non-US cities spanning 3 continents: Europe (London, Stockholm, Berlin, Zurich), Asia (Tokyo), and Australia (Sydney). Fig. 8 shows example data pairs. Furthermore, we leveraged this data to evaluate the generalizability of existing methods and Mapper. Results in Tab. 3 and Fig. 9 show that our model Mapper consistently provides more accurate and realistic zero-shot predictions compared to baselines trained on conventional datasets.
@@ 7 Conclusion
In this work, we propose MIA, a data curation pipeline aimed at empowering anywhere BEV map prediction from FPV images. We release a large MIA dataset obtained through the data engine giving the research community access to $\sim$1.2M FPV-BEV pairs to accelerate anywhere map prediction research. Results from training on the dataset show impressive generalization performance across conventional map prediction datasets, while on the other hand, provides challenging test cases for the research community. Our approach departs from the traditional and expensive autonomous vehicle data collection and labeling paradigm, towards automatic curation of readily-available crowd-sourced data. We believe this work seeds the first step towards anywhere map prediction.
@@ 7.1 Limitations, Biases, Social Impact
While we show promising generalization performance on conventional datasets, we note that label noise inherently exists, to a higher degree than manually collected data, in crowd-sourced data, in both pose correspondence and in BEV map labeling. Such noise is common across large-scale automatically scraped/curated benchmarks such as ImageNet [11]. Moreover, our approach does not capture dynamic classes as they do not exist in static maps. However, we see our approach as indispensable for scale and diversity yet complementary to conventionally obtained datasets.
@@
We thank Mihir Sharma for his work on related research curation and exploration into potential extensions involving terrain height extraction. We thank Mononito Goswami for invaluable discussions on paper organization, framing, experimental design, and project feedback. Additionally, we appreciate David Fan, Simon Stepputtis, and Yaqi Xie for their insightful feedback throughout the project.We are grateful to our reviewers for their constructive feedback.
@@ A.3 Intra-Dataset Class Mappings
A.4 Additional Comparisons between MapItAnywhere and OrienterNet [33]
Figure 11: Qualitative comparisons between different rendering/rasterizing pipelines. ON: OrienterNet[33]. MIA’s rendering pipeline produces semantic map suitable for map prediction.
Figure 12: Detailed comparison of BEV rendering/rasterizing pipelines between OrienterNet[33] (left) and Ours (Right). Our road widths are more satellite aligned than OrienterNet’s roads which are single pixel lines. Our crossings are drawn out from sidewalk to sidewalk whereas OrienterNet crossings are single pixel nodes. We are able to infer sidewalks based on road type in accordance with the OSM standards whereas OrienterNet does not.
MIA  builds on OrienterNet [33] which uses Mapillary [7] and OpenStreetMap [9], to perform RGB image localization within a pre-existing map. Our work extends OrienterNet [33] by using the same data sources to achieve the goal of generalizable FPV-to-BEV map prediction. Our work differs in 2 ways: (1) dataset processing pipeline and (2) released datasets. First, OrienterNet has a fixed set of image IDs that was curated either manually or through a closed-source method. In contrast, we developed a filtering pipeline that allows the automatic curation and collection of images from any location in the world. Users can define a city and change filtering thresholds to meet their needs. Secondly, while OrienterNet-rendered maps provide useful context to an image encoder, they are unsuitable as a prediction target for BEV prediction. This is because the map elements are not consistent with the real world / FPV image. E.g., roads in OrienterNet are 1 pixel wide, whereas the road in ours is a rasterized shape aligned with satellite and FPV images. We illustrate the differences in Figures 11 and 12.
A.5 Additional MIA data samples
Figure 13: Additional MIA data samples and associated metadata
A.6 Mapper Model Architecture
Figure 14: Mapper Architecture
A.7 Hyperparameters
@@
A.8 Dataset Privacy and Consent